A bus travels 225km in 5hrs. Find the average speed in m/s
1
Expert's answer
2019-10-29T10:49:02-0400

1 km=1000 m

225 km= 225000 m

1 hr=60 min

1 min=60 s

1 hr=60*60=3600 s

5 hrs=5*3600= 18000 s

Rate=distance/time=225000 m/18000 s= 12.5 m/s